OpenStudy (anonymous):

A pulley of radius 10 cm turns at 6 revolutions per second. What is the linear velocity of the belt driving the pulley in meters per second? possible answers: 376.99 m/s 1.67 m/s 166.67 m/s 3.77 m/s

OpenStudy (anonymous):

what have you tried so far?

OpenStudy (anonymous):

I just don't know what method to use to do this problem :l

OpenStudy (anonymous):

angular velocity is related to linear velocity by the equation v = wr where v = linear velocity, w = angular velocity, and r = radius first determine the angular velocity (how many radians does the pully travel in one second)? hint: there are 2 pi radians in one rotation

OpenStudy (anonymous):

An alternative (and more direct) method would be to consider exactly how many centimeters 6 rotations is.

OpenStudy (anonymous):

Would the angular velocity be 276.001 rotations per second?

OpenStudy (anonymous):

376*

OpenStudy (anonymous):

Rotations?

jimthompson5910 (jim_thompson5910):

the angular velocity is given at "6 revolutions per second"

OpenStudy (anonymous):

376.991* ugh stupid keyboard

OpenStudy (anonymous):

So then that the answer I got would be the linear velocity?

jimthompson5910 (jim_thompson5910):

The circumference of this circle is C = 2*pi*r C = 2*pi*0.1 ... note: use 0.1 for r because 0.1 meters = 10 cm C = 0.6283185 So if this thing is rotating 6 full rotations a second, then it's covering a linear distance of 6*0.6283185 = 3.769911 meters every second

OpenStudy (anonymous):

btw I meant revolutions not rotations >_<

OpenStudy (anonymous):

Okai awesome that's what I got n_n

jimthompson5910 (jim_thompson5910):

I guess the best way is to imagine this thing is a wheel and you have covered the wheel with paint then you roll the wheel out at 6 rev/sec and it will cover a linear distance of 3.769911 meters

OpenStudy (anonymous):